How do i graph y=2x-1
Answer by math_tutor2020(3835)


This is a linear equation of the form y = mx+b
m = 2 = slope
b = -1 = y intercept

The y intercept is where the line crosses the y axis.
Therefore (0,-1) is one point on the line.

The slope 2 = 2/1 means "go up 2, then right 1".
Follow this motion path to go from (0,-1) to (1,1)

Plot the two points (0,-1) and (1,1)
Draw a straight line through them


Another approach is to plug in x = 0 and it leads to y = -1 through this scratch work
y = 2x-1
y = 2*0-1
y = -1
This shows (0,-1) is on the line
And do the same for x = 1
y = 2x-1
y = 2*1-1
y = 1
This shows (1,1) is also on the line

GeoGebra and Desmos are two (of many more) graphing tools that can be used to verify the answer.